Use the trigonometric substitution to write the algebraic equation as a trigonometric equation of θ, where −π/2 < θ < π/2.
(4 = sqrt (64 − x^2)) , x = 8 sin θ
4=?
Find sin θ and cos θ.
sin θ= ?
cos θ=?

Answer by stanbon(75887)

Use the trigonometric substitution to write the algebraic equation as a trigonometric equation of θ, where −π/2 < θ < π/2.
(4 = sqrt (64 − x^2)) , x = 8 sin θ
16 = 64-x^2
x^2 = 32
x = 4sqrt(2) or x = -4sqrt(2)
-----
Substitute for "x":
8sin(theta) = 4sqrt(2) or 8sin(theta) = -4sqrt(2)
---------------------
sin(theta) = sqrt(2)/2; and cos(theta) = sqrt(2/2) in QI
or sin(theta) = -sqrt(2)/2 and cos(theta) = sqrt(2)/2) in QIV
======================
Find sin θ and cos θ.
sin θ= ?
cos θ=?
======
Cheers,
Stan H.